SpaceRef

SpaceRef


SOHO Daily Report 11 Dec 2002

Status Report From: Goddard Space Flight Center
Posted: Wednesday, December 11, 2002

SOHO DAILY REPORT
Operations Day - 345/2002 (11 December 02)
Mission Day 2567
SOHO FLIGHT OPERATIONS TEAM, NASA/GSFC
Ref. #:  SH-FOT-DR-021211
====================================================================
PASS #1	Madrid			DSS66		0601-1020
PASS #2	Goldstone		DSS27		1324-1615
PASS #3	Goldstone		DSS27		2113-2240
*********************************************************************


PASS #1	Madrid			DSS66		0601-1020

Anomalies:

G02-0658  0624  Lost CMD ability for five minutes due to a PA 
failure.  DR #M101207.
G02-0659  0623  Invalid time-stamp on the data from 0623 until 0638 
due to a TCA problem.
                 DR #M101206.

NRT:

None

Activities Summary:

- SSR Dump
- Collected MDI-M (0616-0634, 0917-0942)
- Collected MDI-H (0942-1001)
- OBT Distribution to SWAN, CELIAS
- SSR SEF Reset (x2)
- SSR Record
------------------


PASS #2	Goldstone		DSS27		1324-1615

Anomalies:

None

NRT:

None

Activities Summary:

- SSR Dump
- Collected MDI-M (1454-1552)
- Uplinked delayed load VIRA21212
- Reaction Wheel Speed Monitoring Daily Adjustment
- CRS Drift Estimation Update
- SSR SEF Reset
- SSR Record
------------------


PASS #3	Goldstone		DSS27		2113-2240

Anomalies:

G02-0660  2011  SH4WS2 and SH4XT1 hung, cause unknown.  Operations 
failed over to String 5
                 until String 4 operational again.
G02-0661  2213  Received degraded TLM from 2213 until 2240.  Lost 
data is unrecoverable.
                 Bad data impacted ability to command, delaying pass 
activities.  Lost 2224
                 magnetogram.  DR #G102005.

NRT:

2120-2210

Activities Summary:

- SSR Dump
- Collected MDI-M (2222-2238)
- Uplinked load TTVC3_IDREC_02345
- SSR SEF Reset
- SSR Record
------------------


-- End of Report --

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.